Science Report – May 07th

Science Report 7 May 2018

Crew 212 – LATAM III

Submitted by GreenHab Officer Hector Palomeque

1. Solar Observation/Astrophotography/Stellar Clusters: The procedure of turning on the MUSK telescope and the dome has been successfully tested. Observations of a galaxy were made with the deep sky objects robotic telescope.

2. X-5: Drones were disassembling for transportation to MDRS: 2 on 3 drones already assembled. Calibration performed. Manual test of 2 on 3 drones scheduled for SOL-3.

3. VESTA: Performing a repair in the landing gear of the drone. First flight test scheduled for May, 09 (SOL 4).

4. Observation Rover: The EVA for today May, 7 (SOL 2) was canceled due to the weather.. A part of the Rover got missing in transit from Grand Junction to MDRS. Critical Component. Solution needed ASAP: Stablish contact with UK.

5. Space Farming: Experiment in standby due to the lack of an operational autoclave. Solution: Dr. Shannon Rupert has explained how to sterilize using a hotplate, alcohol, and bleach.

6. Keep Talking: First round already performed. Subsequent rounds already scheduled.

7. Ethnography study: First observation started yesterday (SOL-1). Waiting for sending audio-information to receive the feedback from the Colombian team project.

8. Pulse Oximeter Measurement: First measurement made yesterday. Data recorded.

9. Software for macroscopic characterization: Code for detecting counts in progress. There is an issue with detecting the petri dish. Solution: process more photographs needed (soil bacteria growing in plates).

Copyright © The Mars Society. All rights reserved. | Main Site